### v1

# VER-Bench: Evaluating MLLMs on Reasoning with Fine-Grained Visual Evidence

Abstract.
With the rapid development of MLLMs, evaluating their visual capabilities has become increasingly crucial. Current benchmarks primarily fall into two main types: basic perception benchmarks, which focus on local details but lack deep reasoning (e.g., ”what is in the image?”), and mainstream reasoning benchmarks, which concentrate on prominent image elements but may fail to assess subtle clues requiring intricate analysis. However, profound visual understanding and complex reasoning depend more on interpreting subtle, inconspicuous local details than on perceiving salient, macro-level objects. These details, though occupying minimal image area, often contain richer, more critical information for robust analysis. To bridge this gap, we introduce the VER-Bench, a novel framework to evaluate MLLMs’ ability to: 1) identify fine-grained visual clues, often occupying on average just 0.25% of the image area; 2) integrate these clues with world knowledge for complex reasoning. Comprising 374 carefully designed questions across Geospatial, Temporal, Situational, Intent, System State, and Symbolic reasoning, each question in VER-Bench is accompanied by structured evidence: visual clues and question-related reasoning derived from them. VER-Bench reveals current models’ limitations in extracting subtle visual evidence and constructing evidence-based arguments, highlighting the need to enhance models’ capabilities in fine-grained visual evidence extraction, integration, and reasoning for genuine visual understanding and human-like analysis. Dataset and additional materials are available https://github.com/verbta/ACMMM-25-Materials.
1. Introduction
The rapid evolution of multimodal large language models (MLLMs) has unlocked impressive capabilities in understanding and generating content that seamlessly blends vision and language (Wu et al., 2024; Wang et al., 2024b; Li et al., 2024a). These models can now describe images with remarkable accuracy (Bucciarelli et al., 2025; Kim et al., 2025), answer visual questions with nuanced responses (Hu et al., 2024; Jian et al., 2024), and even generate images based on textual prompts (Song et al., 2024; Feng et al., 2025). However, a critical frontier remains largely unexplored: assessing whether these models genuinely reason by meticulously extracting and integrating specific visual evidence, or if their inferences sometimes rely more on learned textual priors or a more holistic understanding of obvious visual features. This issue is particularly pertinent when tackling complex inferences that require delving into subtle details, rather than merely leveraging high-level visual clues.
While several benchmarks now evaluate the visual capabilities of MLLMs, their focus often deviates from the specific demands of rigorous, locally evidence-based visual reasoning that VER-Bench aims to assess. As illustrated in the left sub-figure of Fig.1, certain benchmarks designed for basic visual perception, such as VStar (Wu and Xie, 2024) and MME-RealWorld (Fu et al., 2024), primarily emphasize identifying ”what is in the image,” focusing on local, minute details. However, they typically lack tasks that require deeper inferential reasoning grounded in those details. On the other hand, benchmarks incorporating Chain-of-Thought (CoT) methodologies, such as MME-CoT (Jiang et al., 2025) and Visual CoT (Shao et al., 2024), aim to model reasoning processes. The visual grounding for these reasoning steps often relies on salient or holistic aspects of the image, rather than explicitly identifying subtle, fine-grained visual evidence.
This motivates our focus on small image regions, which often provide more decisive information than larger, salient areas. For example, in the right sub-figure of Fig.1, the distant statue in the bay serves as a distinctive landmark for precise localization, whereas the prominent skyline and building clusters offer limited diagnostic value. Similarly, while a cruise ship’s size or color provides ambiguous cues, fine details such as a hull number or unique logo enable definitive identification
To bridge this evaluation gap, we introduce VER-Bench, the Visual Evidence-based Reasoning Benchmark. This benchmark is specifically designed to assess an MLLM’s ability to first extract minute yet crucial visual clues from images, and then construct a logical reasoning chain directly grounded in this granular evidence to support its conclusions. VER-Bench encompasses six challenging reasoning domains, each crafted to mirror diverse real-world analytical tasks:
- Geospatial Reasoning: Inferring location based on subtle geographic or cultural indicators.
- Temporal Reasoning: Determining time periods or sequences from visual context.
- Situational Adaptation Reasoning: Assessing appropriate actions or understanding evolving scenarios.
- Intent Inference: Deducing the purpose or goals of agents within the scene.
- System State Understanding: Analyzing the operational status or condition of depicted systems.
- Symbolic Analysis: Interpreting abstract symbols, codes, or non-linguistic visual information.
Each instance in VER-Bench is meticulously annotated with detailed visual evidence pairs, linking specific, often small and localized visual clues directly to the reasoning steps they support. This design mirrors sophisticated human reasoning, demanding that models identify subtle visual details—such as faint architectural patterns or small text—and construct verifiable, evidence-based reasoning chains. Each question is structured around a reasoning chain that includes: (1) Precise Visual Clues (e.g., small textual signs, symbols, minute environmental details) and (2) Step-by-Step Reasoning that logically connects these clues to the conclusion. The primary contributions of this work are summarized as follows:
Conceptualization of Visual Evidence-Based Reasoning: We formalize visual reasoning grounded in fine-grained evidence, highlighting its importance for rigorous MLLM evaluation.
Introduction of VER-Bench: We introduce a novel benchmark with tailored metrics to assess both final answer accuracy and the fidelity of evidence/reasoning chains.
Building a Scalable Data Pipeline: We detail a systematic pipeline for generating high-quality, diverse datasets to support scalable benchmark expansion.
2. Related Work
2.1. Multimodal Large Language Models
The development of multimodal large language models (MLLMs) (Liu et al., 2023b; Zhu et al., 2023; Lin et al., 2023; Wang et al., 2024a) represents a significant advancement in AI, building upon Large Language Models (LLMs) (Touvron et al., 2023; Team, 2024) and vision models (Radford et al., 2021). While closed-source models like GPT-4o (OpenAI, 2024a) demonstrate strong capabilities, the research community has actively developed open-source alternatives such as LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), InternVL2 (Chen et al., 2024c), and Qwen2-VL (Wang et al., 2024a). These advancements are driven by improved multimodal data (Chen et al., 2025; Liu et al., 2024a; Wang et al., 2023b; Ye et al., 2023), novel alignment techniques (Bai et al., 2023; Dong et al., 2024; Liu et al., 2024b; Li et al., 2024b; Wang et al., 2023a), and methods like LoRA (Hu et al., 2021). Recent research has focused on enhancing reasoning capabilities, with models like o1 (OpenAI, 2024b) and works on multi-step Chain-of-Thought (CoT) reasoning with self-reflection (Team, 2024; Du et al., 2025). Despite increased accessibility through commercial APIs such as GPT-4v (Achiam et al., 2023) and Gemini-Pro-V (Team et al., 2023), questions regarding real-world efficacy remain (Zhang et al., 2024; Fu et al., 2023).
2.2. MLLM Benchmarks
The rapid advancement of MLLMs has spurred the creation of comprehensive evaluation benchmarks (Duan et al., 2024; Fu et al., 2024). These benchmarks cover a wide range of capabilities, including real-world applications and visual perception (e.g., MME (Fu et al., 2024), MMBench (Liu et al., 2023a), MMVet (Yu et al., 2023)), integrated reasoning (e.g., MMStar (Chen et al., 2024a), MMMU (Yue et al., 2024)), and video understanding (e.g., MMBench-Video (Fang et al., 2025)). Specialized benchmarks also exist for tasks like chart interpretation (ChartQA (Masry et al., 2022)), document analysis (DocVQA (Mathew et al., 2021)), and mathematical reasoning (Yan et al., 2024). Evaluation methodologies primarily include Visual Question Answering (VQA), which allows open-ended responses but faces scoring objectivity issues (Papineni et al., 2002; Vedantam et al., 2015; Anderson et al., 2016), and Multiple Choice Questions (MCQ) (Schwenk et al., 2022), which are prevalent in leaderboards due to more reliable outcomes. However, existing MCQ benchmarks face several criticisms. Older benchmarks like MME (Fu et al., 2024) and POPE (Li et al., 2023) are often saturated, while newer, more challenging ones (e.g., MME-RealWorld (Zhang et al., 2024), RealWorldQA (XAI, 2024)) can suffer from ambiguity. Other concerns include answer leakage, partly addressed since MMStar (Chen et al., 2024a), and data redundancy or overlap (Zhang et al., 2025). Additionally, there is a recognized deficiency in benchmarks for specific industrial applications and non-English languages, with limited options like MME-Realworld-CN (Zhang et al., 2024) and MMBench-CN (Liu et al., 2023a).
3. Dateset Construction
The images in VER-Bench were sourced from the internet, primarily depicting real-world scenes. Over 3000 images are collected, with an average resolution of around 15002000 pixels.
3.1. Annotation Pipeline and Curation Process
The dataset is built via a multi-stage annotation pipeline with four key steps (as shown in Fig. 2) to generate high-quality visual evidence reasoning chains for real-world scenarios.
Step 1: Initial Annotation. With image as visual input, Gemini-2.5-Pro model is then tasked to: 1. Analyze visual details from the image. 2. Construct a realistic scenario based on the observation. 3. Design a context-driven reasoning question. 4. Provide logical answers supported by visual details.
Step 2: Scenario and Question Assessment. To ensure alignment between scenarios and visual content, and the relevance of questions, we employ two validation procedures: 1. GPT-4o evaluates the relevance between the masked image (randomly masked by 50% area) and the scenario; 2. Multiple random subsets are analyzed using GPT-4o relevance metrics. Entries with lower original image-question relevance compared to alternative pairings are removed. These strategies ensure that only image-scenario-question pairs with strong mutual relevance are retained.
Step 3: Answer and Visual Evidence Verification. This step ensured the dataset’s internal consistency and factual accuracy. We cross-verified answers, reasoning, and supporting visual evidence using multiple MLLMs (e.g., Qwen2.5-VL-72B and Claude-3.5-Sonnet). Only instances with consensus on the final answer and reasoning chain across all models were retained, minimizing model-specific biases and enhancing annotation reliability.
Step 4: Human Expert Review. Human experts conducted two review rounds to ensure accuracy and remove errors: 1. Visual Clue Verification: Experts confirmed the presence of each annotated visual clue in the image, eliminating any incorrect or hallucinated clues; 2. Reasoning Validation: Each reasoning step was reviewed for logical validity and alignment with real-world knowledge.
Additionally, bounding boxes were manually annotated for all visual clues to precisely locate them within the image, enhancing the reasoning chain’s interpretability and traceability.
3.2. Dataset Overview
The dataset comprises 342 images and 374 Q&A pairs, with 614 visual clues annotated by 977 bounding boxes. These clues cover an average of 0.25% of the image area, making them ideal for fine-grained reasoning evaluation. VER-Bench is categorized into six tasks (Fig. 3), detailed below:
Geospatial Reasoning. Identifies geographic locations using visual clues like landmarks, scripts, symbols, and vehicle details. It requires global geography knowledge and cross-referencing multiple clues for accurate localization.
Temporal Reasoning. Infers time information (e.g., dates) from clues like clocks, text, seasonal signs, and event-specific elements.
Situational Adaptation Reasoning. Interprets complex scenes to infer situational contexts (e.g., events, activities) by understanding human behavior, spatial use, and environmental cues.
Intent Inference. Uncovers the underlying purpose of objects, designs, or behaviors, requiring semantic abstraction and contextual interpretation.
System State Understanding. Assesses the operational status of facilities or objects by analyzing visual signs like warning labels, crowd flow, or device status.
Symbolic Analysis. Recognizes and interprets cultural or historical symbols (e.g., emblems, icons) using both visual perception and deep cultural knowledge.
3.3. Evaluation Metric of VER-Bench
VER-Bench evaluates models using four metrics: Answer Correctness, Clue Coverage, Reasoning Consistency, and Evidence-Answer Relevance. During evaluation, models are prompted to follow a structured output format (details in Appendix):
where represents the final answer, denotes the -th visual clue identified by the model, and is the reasoning based on the clue.
The corresponding ground truth is defined as:
where denotes the correct answer, and , represent the -th visual clue and its corresponding reasoning from the ground-truth (GT). Each instance is then scored along the following axes:
Answer Correctness (AC): The AC is determined by the semantic similarity between the model’s answer and the GT answer . The higher , the greater alignment in meaning and conclusion.
Clue Coverage (CC): This metric evaluates how completely the model captures the information in the ground truth. The higher indicates that more critical visual clues are correctly recognized and described by the model.
where , .
Reasoning Quality (RQ): This metric measures how well the model’s reasoning consistent with the ground-truth reasoning. A high implies that the model’s logic is semantically aligned with the expected reasoning process.
where , .
Evidence-Answer Relevance (ER): The measure whether each reasoning based on visual clue leads to the final answer and whether any reasoning is off-topic or irrelevant.
The LLM used to score answer correctness, clue coverage, reasoning quality, and evidence-answer relevance is GPT-4. All scores range from 0 to 10. The final average accuracy is computed as:
where is the total number of instances in VER-Bench.
3.4. Holistic Evidence Evaluation vs. Individual Clue-Reasoning Matching
In and metrics, we do not first use LLM to judge whether each model output evidence (clue-reasoning pair) matches the GT evidence and then calculate recall based on matches. Instead, we concatenate all model output clues and GT clues separately and directly evaluate their overall matching score using LLM.
Visual reasoning differs from problems requiring strict sequential logic (e.g., math, programming). In visual reasoning, clue-reasoning pairs often lack strong sequential dependencies, and the order of identification typically does not affect the outcome. These pairs act as complementary ”evidence,” with more identified pairs enhancing answer reliability. Additionally, imperfect alignment between model-output and GT clues results in non-binary matching, complicating recall calculations.
4. Experiments and Analysis
In this section, we systematically evaluate state-of-the-art (SOTA) models on VER-Bench. We detail the experiment setup in Sec. 4.1 and report quantitative results and analysis insights in Sec. 4.2.
4.1. Experiment Setups
We evaluated 12 open-source models, including Qwen2.5-VL (8B, 32B, 72B) (Bai et al., 2023), Qwen2-VL (8B) (Wang et al., 2024a), InternVL3 (8B, 38B, 78B) (Zhu et al., 2025), InternVL2.5 (8B) (Chen et al., 2024b), InternVL2 (8B) (Dong et al., 2024), MiniCPM-V-2.6 and MiniCPM-o-2.6 (Yao et al., 2024), and LLaVA-OneVision-72B (Li et al., 2024c). Additionally, we evaluated three closed-source models: Gemini-2.5-Pro-Preview (Team et al., 2023), GPT-4o (Achiam et al., 2023), and Claude-3.7-Sonnet (Anthropic, [n. d.]). Open-source models are tested on NVIDIA H20 GPUs using the MS-Swift framework, while closed-source models are evaluated via API calls. GPT-4o was used to score predictions against ground truth.
Average Accuracy | ||||||||||
| Open-source MLLMs | ||||||||||||||||
| MiniCPM-o 2.6 | Qwen2.5-7B (Yang et al., 2025b) | 42.4 | 20.5 | 28.8 | 76.0 | 41.9 | ||||||||||
| MiniCPM-V 2.6 | Qwen2-7B (Yang et al., 2025a) | 41.0 | 20.2 | 27.3 | 82.1 | 42.7 | ||||||||||
| LLaVA-OneVision-72B | Qwen2-72B (Yang et al., 2025a) | 48.8 | 24.7 | 33.7 | 86.5 | 48.4 | ||||||||||
| InternVL2-8B | InternLM-7B (Cai et al., 2024) | 35.2 | 16.5 | 22.2 | 83.9 | 39.4 | ||||||||||
| InternVL2.5-8B | InternLM-7B (Cai et al., 2024) | 38.9 | 19.8 | 25.5 | 77.5 | 40.4 | ||||||||||
| InternVL2.5-38B | Qwen2.5-32B (Yang et al., 2025b) | 46.6 | 30.0 | 35.3 | 87.4 | 49.8 | ||||||||||
| InternVL3-8B | Qwen2.5-7B (Yang et al., 2025b) | 47.4 | 31.8 | 39.0 | 83.0 | 50.3 | ||||||||||
| InternVL3-78B | Qwen2.5-72B (Yang et al., 2025b) | 55.1 | 40.6 | 45.7 | 86.2 | 56.9 | ||||||||||
| Qwen2-VL-7B | Qwen2-7B (Yang et al., 2025a) | 52.2 | 26.6 | 34.3 | 79.8 | 48.2 | ||||||||||
| Qwen2.5-VL-7B | Qwen2.5-7B (Yang et al., 2025b) | 52.9 | 29.7 | 34.4 | 84.8 | 50.4 | ||||||||||
| Qwen2.5-VL-72B | Qwen2.5-72B (Yang et al., 2025b) | 61.8 | 38.2 | 46.6 | 92.0 | 59.6 | ||||||||||
| Qwen2.5-VL-32B | Qwen2.5-32B (Yang et al., 2025b) | 63.6 | 49.6 | 45.6 | 84.4 | 60.8 | ||||||||||
| Closed-source MLLMs | ||||||||||||||||
| Claude-3.7-Sonnet-20250219 |
52.7 | 37.2 | 43.4 | 92.1 | 56.3 | ||||||||||
| GPT-4o |
62.8 | 46.7 | 54.5 | 93.3 | 64.4 | ||||||||||
| Gemini-2.5-Pro-Preview-05-06 |
79.0 | 67.9 | 69.4 | 90.7 | 76.8 | ||||||||||
4.2. Quantitative Results
Table 1 evaluates open-source and closed-source MLLMs on VER Bench, using four metrics: Answer Correctness (AC), Clue Coverage (CC), Reasoning Quality (RQ), and Evidence-Answer Relevance (ER). These metrics collectively assess not only the final answer accuracy but also the model’s ability to identify relevant visual clues, perform logical reasoning, and align evidence with responses.
Closed-source models generally outperform open-source ones, with Gemini-2.5-Pro-Preview achieving the highest Average Accuracy of 76.8, surpassing GPT-4o by 12.5 points. Among open-source models, Qwen2.5-VL-32B achieved the top score of 60.8, outperforming larger models like Qwen2.5-VL-72B and InternVL3-78B. This is attributed to its reinforcement learning strategy, which enhances fine-grained visual analysis and instruction following. Qwen2.5-VL-72B closely follows at 59.6, while InternVL3-78B scored 56.9. LLaVA-OneVision-72B, despite its larger parameters, performs comparably to smaller Qwen-VL and InternVL models, likely due to its focus on cross-modal migration to multi-images and videos rather than single-image understanding. The edge models MiniCPM-V 2.6 and MiniCPM-o 2.6 show strong performance for their size, surpassing InternVL2-8B and InternVL2.5-8B.
4.2.1. Answer Correctness
This metric assesses the semantic similarity between the model-generated final answer and the GT answer. Higher scores indicate greater alignment in meaning and conclusion. Gemini-2.5-Pro-Preview achieves the highest , demonstrating its proficiency in generating semantically accurate answers. Qwen2.5-VL-32B (61.8) and GPT-4o (62.8) also show strong performance. Conversely, models like InternVL2-8B (35.2) have significantly lower , indicating difficulties in reaching the correct conclusion despite potential strengths in other aspects (e.g., clue identification or reasoning). This metric is crucial as it directly reflects the model’s ability to fulfill the primary task objective.
4.2.2. Clue Coverage
Clue Coverage () measures how well models identify GT visual clues. Higher indicates better recognition of critical visual information, leading to more accurate localization of relevant image regions and higher answer accuracy. For example, Qwen-VL-series models saw increase from 26.6 (Qwen2-VL-7B) to 38.2 (Qwen2.5-VL-72B), correlating with improved answer correctness. However, many lower-performing open-source models (e.g., InternVL2-8B with 16.5 ) struggle significantly, often generating excessive irrelevant clues. This redundancy disrupts the correlation assessment with GT clues, likely due to poor instruction-following. Low suggests models overlook essential visual details, impairing reasoning and answer correctness, thus highlighting the importance of effective visual perception in MLLMs.
4.2.3. Reasoning Quality
measures the semantic consistency between the reasoning generated by the model and the GT reasoning. A high value signifies that the model’s logical deduction process closely aligns with the expected correct reasoning process. Gemini-2.5-Pro-Preview leads with 69.4 points, demonstrating its ability not only to identify clues correctly but also to explain how these clues lead to conclusions in a manner akin to human standards. GPT-4o (54.5) and Claude-3.7-Sonnet (43.4) also exhibit strong reasoning quality. Among open-source models, Qwen2.5-VL-72B (46.6), InternVL3-78B (45.7), and Qwen2.5-VL-32B (45.6) stand out, indicating they excel at simulating correct reasoning paths compared to other open-source counterparts. In contrast, models such as InternVL2.5-8B (25.5) and InternVL2-8B (22.2) score lower. This suggests their reasoning processes may contain logical flaws, insufficient explanations, or significant semantic deviations from GT reasoning—even if they identify some correct visual clues.
4.2.4. Evidence-to-Answer Relevance
Across all metrics, the average score for is significantly higher than others, with even the lowest score exceeding 76. This indicates that even if a model’s final answer () or reasoning process () deviates from the GT, it can still construct internally self-consistent ”evidence-reasoning-answer”. For instance, Qwen2.5-VL-72B outperformed closed-source models like Gemini-2.5-Pro-Preview in , highlighting its potential for generating coherent explanations. Conversely, lower scores, such as MiniCPM-o 2.6 (76.0) and InternVL2.5-8B (77.5), suggest that these models’ reasoning steps may be more divergent or contain redundant information with limited support for their conclusions. is crucial for evaluating the coherence and persuasiveness of model outputs. A high score signifies a more rigorous and focused argumentation process.
and together show how well a model can think and if its thinking is consistent. High means the model can copy good thinking. High means the model can organize its thoughts to back up its conclusions, even if they aren’t perfect. Gemini-2.5-Pro-Preview-05-06 excels in both metrics, demonstrating strong logical reasoning and consistency. In contrast, open-source models generally perform adequately in but exhibit a significant gap in compared to top closed-source models. This discrepancy suggests that future research should focus on enhancing the logical coherence of reasoning in open-source models, bringing them closer to human-like reasoning patterns. Overall, these two metrics provide valuable insights into the reasoning capabilities and limitations of MLLMs when tackling complex tasks.
5. Conclusion
In this paper, we introduce VER-Bench, a benchmark tailored to evaluate the visual reasoning capabilities of MLLMs. It challenges models to identify critical visual details, interpret their meanings, and integrate fragmented information to derive reliable answers. This process mirrors human visual analysis, where identifying abundant supporting evidence is crucial for strengthening conclusions. Our evaluation results demonstrate a positive correlation between Clue Coverage and Answer Correctness, validating our motivation. VER-Bench represents a significant advancement in the evaluation and development of MLLMs. By promoting nuanced, evidence-driven reasoning, it encourages models to identify and integrate fine-grained visual clues, thereby forming robust conclusions. This framework not only pushes models beyond surface-level understanding but also drives the development of more sophisticated reasoning capabilities. Through VER-Bench, we aim to foster the evolution of MLLMs that approach visual reasoning tasks with the depth and precision characteristic of human cognition.
Acknowledgements.
Acknowledgements: This work was supported in part by the Key Deployment Program of the Chinese Academy of Sciences, China under Grant KGFZD145-25-39 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant E1XA310103.References
- (1)
- Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774 (2023).
- Anderson et al. (2016) Peter Anderson, Basura Fernando, Mark Johnson, and Stephen Gould. 2016. Spice: Semantic propositional image caption evaluation. In Computer Vision–ECCV 2016: 14th European Conference, Amsterdam, The Netherlands, October 11-14, 2016, Proceedings, Part V 14. Springer, 382–398.
- Anthropic ([n. d.]) S Anthropic. [n. d.]. Model card addendum: Claude 3.5 haiku and upgraded claude 3.5 sonnet. URL https://api. semanticscholar. org/CorpusID 273639283 ([n. d.]).
- Bai et al. (2023) Jinze Bai, Shuai Bai, Shusheng Yang, Shijie Wang, Sinan Tan, Peng Wang, Junyang Lin, Chang Zhou, and Jingren Zhou. 2023. Qwen-vl: A frontier large vision-language model with versatile abilities. arXiv preprint arXiv:2308.12966 (2023).
- Bucciarelli et al. (2025) Davide Bucciarelli, Nicholas Moratelli, Marcella Cornia, Lorenzo Baraldi, and Rita Cucchiara. 2025. Personalizing multimodal large language models for image captioning: an Experimental analysis. In European Conference on Computer Vision. Springer, 351–368.
- Cai et al. (2024) Zheng Cai, Maosong Cao, Haojiong Chen, Kai Chen, Keyu Chen, Xin Chen, Xun Chen, Zehui Chen, Zhi Chen, Pei Chu, et al. 2024. Internlm2 technical report. arXiv preprint arXiv:2403.17297 (2024).
- Chen et al. (2025) Lin Chen, Jinsong Li, Xiaoyi Dong, Pan Zhang, Conghui He, Jiaqi Wang, Feng Zhao, and Dahua Lin. 2025. Sharegpt4v: Improving large multi-modal models with better captions. In European Conference on Computer Vision. Springer, 370–387.
- Chen et al. (2024a) Lin Chen, Jinsong Li, Xiaoyi Dong, Pan Zhang, Yuhang Zang, Zehui Chen, Haodong Duan, Jiaqi Wang, Yu Qiao, Dahua Lin, and Feng Zhao. 2024a. Are We on the Right Way for Evaluating Large Vision-Language Models? arXiv preprint arXiv:2403.20330 (2024).
- Chen et al. (2024b) Zhe Chen, Weiyun Wang, Yue Cao, Yangzhou Liu, Zhangwei Gao, Erfei Cui, Jinguo Zhu, Shenglong Ye, Hao Tian, Zhaoyang Liu, et al. 2024b. Expanding performance boundaries of open-source multimodal models with model, data, and test-time scaling. arXiv preprint arXiv:2412.05271 (2024).
- Chen et al. (2024c) Zhe Chen, Weiyun Wang, Hao Tian, Shenglong Ye, Zhangwei Gao, Erfei Cui, Wenwen Tong, Kongzhi Hu, Jiapeng Luo, Zheng Ma, et al. 2024c. How far are we to gpt-4v? closing the gap to commercial multimodal models with open-source suites. arXiv preprint arXiv:2404.16821 (2024).
- Dong et al. (2024) Xiaoyi Dong, Pan Zhang, Yuhang Zang, Yuhang Cao, Bin Wang, Linke Ouyang, Xilin Wei, Songyang Zhang, Haodong Duan, Maosong Cao, et al. 2024. Internlm-xcomposer2: Mastering free-form text-image composition and comprehension in vision-language large model. arXiv preprint arXiv:2401.16420 (2024).
- Du et al. (2025) Yifan Du, Zikang Liu, Yifan Li, Wayne Xin Zhao, Yuqi Huo, Bingning Wang, Weipeng Chen, Zheng Liu, Zhongyuan Wang, and Ji-Rong Wen. 2025. Virgo: A Preliminary Exploration on Reproducing o1-like MLLM. arXiv preprint arXiv:2501.01904 (2025).
- Duan et al. (2024) Haodong Duan, Junming Yang, Yuxuan Qiao, Xinyu Fang, Lin Chen, Yuan Liu, Xiaoyi Dong, Yuhang Zang, Pan Zhang, Jiaqi Wang, Dahua Lin, and Kai Chen. 2024. VLMEvalKit: An Open-Source Toolkit for Evaluating Large Multi-Modality Models. arXiv:2407.11691 [cs.CV] https://arxiv.org/abs/2407.11691
- Fang et al. (2025) Xinyu Fang, Kangrui Mao, Haodong Duan, Xiangyu Zhao, Yining Li, Dahua Lin, and Kai Chen. 2025. Mmbench-video: A long-form multi-shot benchmark for holistic video understanding. NeurIPS (2025).
- Feng et al. (2025) Dong Feng, Ping Guo, Encheng Peng, Mingmin Zhu, Wenhao Yu, and Peng Wang. 2025. PoseLLaVA: Pose Centric Multimodal LLM for Fine-Grained 3D Pose Manipulation. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 39. 2951–2959.
- Fu et al. (2023) Chaoyou Fu, Peixian Chen, Yunhang Shen, Yulei Qin, Mengdan Zhang, Xu Lin, Jinrui Yang, Xiawu Zheng, Ke Li, Xing Sun, Yunsheng Wu, and Rongrong Ji. 2023. MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models. arXiv preprint arXiv:2306.13394 (2023).
- Fu et al. (2024) Chaoyou Fu, Yi-Fan Zhang, Shukang Yin, Bo Li, Xinyu Fang, Sirui Zhao, Haodong Duan, Xing Sun, Ziwei Liu, Liang Wang, et al. 2024. Mme-survey: A comprehensive survey on evaluation of multimodal llms. arXiv (2024).
- Hu et al. (2021) Edward J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2021. Lora: Low-rank adaptation of large language models. arXiv preprint arXiv:2106.09685 (2021).
- Hu et al. (2024) Wenbo Hu, Yifan Xu, Yi Li, Weiyue Li, Zeyuan Chen, and Zhuowen Tu. 2024. Bliva: A simple multimodal llm for better handling of text-rich visual questions. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 2256–2264.
- Jian et al. (2024) Pu Jian, Donglei Yu, and Jiajun Zhang. 2024. Large language models know what is key visual entity: An llm-assisted multimodal retrieval for vqa. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing. 10939–10956.
- Jiang et al. (2025) Dongzhi Jiang, Renrui Zhang, Ziyu Guo, Yanwei Li, Yu Qi, Xinyan Chen, Liuhui Wang, Jianhan Jin, Claire Guo, Shen Yan, et al. 2025. MME-CoT: Benchmarking Chain-of-Thought in Large Multimodal Models for Reasoning Quality, Robustness, and Efficiency. arXiv preprint arXiv:2502.09621 (2025).
- Kim et al. (2025) Jaeyoung Kim, Jongho Lee, Hong-Jun Choi, Ting-Yao Hsu, Chieh-Yang Huang, Sungchul Kim, Ryan Rossi, Tong Yu, Clyde Lee Giles, Ting-Hao’Kenneth’ Huang, et al. 2025. Multi-LLM Collaborative Caption Generation in Scientific Documents. arXiv preprint arXiv:2501.02552 (2025).
- Li et al. (2024c) Bo Li, Yuanhan Zhang, Dong Guo, Renrui Zhang, Feng Li, Hao Zhang, Kaichen Zhang, Yanwei Li, Ziwei Liu, and Chunyuan Li. 2024c. LLaVA-OneVision: Easy Visual Task Transfer. arXiv:2408.03326 [cs.CV] https://arxiv.org/abs/2408.03326
- Li et al. (2024a) Wei Li, Hehe Fan, Yongkang Wong, Yi Yang, and Mohan Kankanhalli. 2024a. Improving context understanding in multimodal large language models via multimodal composition learning. In Forty-first International Conference on Machine Learning.
- Li et al. (2023) Yifan Li, Yifan Du, Kun Zhou, Jinpeng Wang, Xin Zhao, and Ji-Rong Wen. 2023. Evaluating Object Hallucination in Large Vision-Language Models. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore.
- Li et al. (2024b) Zhang Li, Biao Yang, Qiang Liu, Zhiyin Ma, Shuo Zhang, Jingxu Yang, Yabo Sun, Yuliang Liu, and Xiang Bai. 2024b. Monkey: Image resolution and text label are important things for large multi-modal models. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 26763–26773.
- Lin et al. (2023) Ziyi Lin, Chris Liu, Renrui Zhang, Peng Gao, Longtian Qiu, Han Xiao, Han Qiu, Chen Lin, Wenqi Shao, Keqin Chen, et al. 2023. Sphinx: The joint mixing of weights, tasks, and visual embeddings for multi-modal large language models. ECCV 2024 (2023).
- Liu et al. (2024b) Dongyang Liu, Renrui Zhang, Longtian Qiu, Siyuan Huang, Weifeng Lin, Shitian Zhao, Shijie Geng, Ziyi Lin, Peng Jin, Kaipeng Zhang, et al. 2024b. Sphinx-x: Scaling data and parameters for a family of multi-modal large language models. arXiv preprint arXiv:2402.05935 (2024).
- Liu et al. (2024a) Haotian Liu, Chunyuan Li, Yuheng Li, and Yong Jae Lee. 2024a. Improved baselines with visual instruction tuning. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 26296–26306.
- Liu et al. (2023b) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2023b. Visual Instruction Tuning. In NeurIPS.
- Liu et al. (2023a) Yuan Liu, Haodong Duan, Yuanhan Zhang, Bo Li, Songyang Zhang, Wangbo Zhao, Yike Yuan, Jiaqi Wang, Conghui He, Ziwei Liu, Kai Chen, and Dahua Lin. 2023a. MMBench: Is Your Multi-modal Model an All-around Player? arXiv preprint arXiv:2307.06281 (2023).
- Masry et al. (2022) Ahmed Masry, Do Xuan Long, Jia Qing Tan, Shafiq Joty, and Enamul Hoque. 2022. Chartqa: A benchmark for question answering about charts with visual and logical reasoning. arXiv preprint arXiv:2203.10244 (2022).
- Mathew et al. (2021) Minesh Mathew, Dimosthenis Karatzas, and CV Jawahar. 2021. Docvqa: A dataset for vqa on document images. In Proceedings of the IEEE/CVF winter conference on applications of computer vision. 2200–2209.
- OpenAI (2024a) OpenAI. 2024a. Hello GPT-4o. https://openai.com/index/hello-gpt-4o/
- OpenAI (2024b) OpenAI. 2024b. Introducing openai o1, 2024. https://openai.com/o1/
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th annual meeting of the Association for Computational Linguistics. 311–318.
- Radford et al. (2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, Gretchen Krueger, and Ilya Sutskever. 2021. Learning Transferable Visual Models From Natural Language Supervision. In International Conference on Machine Learning. https://api.semanticscholar.org/CorpusID:231591445
- Schwenk et al. (2022) Dustin Schwenk, Apoorv Khandelwal, Christopher Clark, Kenneth Marino, and Roozbeh Mottaghi. 2022. A-OKVQA: A Benchmark for Visual Question Answering Using World Knowledge. In Computer Vision – ECCV 2022: 17th European Conference, Tel Aviv, Israel, October 23–27, 2022, Proceedings, Part VIII (Tel Aviv, Israel). Springer-Verlag, Berlin, Heidelberg, 146–162. https://doi.org/10.1007/978-3-031-20074-8_9
- Shao et al. (2024) Hao Shao, Shengju Qian, Han Xiao, Guanglu Song, Zhuofan Zong, Letian Wang, Yu Liu, and Hongsheng Li. 2024. Visual cot: Advancing multi-modal language models with a comprehensive dataset and benchmark for chain-of-thought reasoning. Advances in Neural Information Processing Systems 37 (2024), 8612–8642.
- Song et al. (2024) Kunpeng Song, Yizhe Zhu, Bingchen Liu, Qing Yan, Ahmed Elgammal, and Xiao Yang. 2024. Moma: Multimodal llm adapter for fast personalized image generation. In European Conference on Computer Vision. Springer, 117–132.
- Team et al. (2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, Katie Millican, et al. 2023. Gemini: a family of highly capable multimodal models. arXiv preprint arXiv:2312.11805 (2023).
- Team (2024) Qwen Team. 2024. QVQ: To See the World with Wisdom. https://qwenlm.github.io/blog/qvq-72b-preview/
- Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
- Vedantam et al. (2015) Ramakrishna Vedantam, C Lawrence Zitnick, and Devi Parikh. 2015. Cider: Consensus-based image description evaluation. In Proceedings of the IEEE conference on computer vision and pattern recognition. 4566–4575.
- Wang et al. (2023b) Junke Wang, Lingchen Meng, Zejia Weng, Bo He, Zuxuan Wu, and Yu-Gang Jiang. 2023b. To see is to believe: Prompting gpt-4v for better visual instruction tuning. arXiv preprint arXiv:2311.07574 (2023).
- Wang et al. (2024a) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, et al. 2024a. Qwen2-vl: Enhancing vision-language model’s perception of the world at any resolution. arXiv preprint arXiv:2409.12191 (2024).
- Wang et al. (2023a) Weihan Wang, Qingsong Lv, Wenmeng Yu, Wenyi Hong, Ji Qi, Yan Wang, Junhui Ji, Zhuoyi Yang, Lei Zhao, Xixuan Song, et al. 2023a. Cogvlm: Visual expert for pretrained language models. arXiv preprint arXiv:2311.03079 (2023).
- Wang et al. (2024b) Yabing Wang, Le Wang, Qiang Zhou, Zhibin Wang, Hao Li, Gang Hua, and Wei Tang. 2024b. Multimodal llm enhanced cross-lingual cross-modal retrieval. In Proceedings of the 32nd ACM International Conference on Multimedia. 8296–8305.
- Wu and Xie (2024) Penghao Wu and Saining Xie. 2024. V?: Guided visual search as a core mechanism in multimodal llms. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 13084–13094.
- Wu et al. (2024) Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua. 2024. Next-gpt: Any-to-any multimodal llm. In Forty-first International Conference on Machine Learning.
- XAI (2024) XAI. 2024. grok-1.5v. https://x.ai/blog/grok-1.5v
- Yan et al. (2024) Yibo Yan, Shen Wang, Jiahao Huo, Hang Li, Boyan Li, Jiamin Su, Xiong Gao, Yi-Fan Zhang, Tianlong Xu, Zhendong Chu, et al. 2024. Errorradar: Benchmarking complex mathematical reasoning of multimodal large language models via error detection. arXiv (2024).
- Yang et al. (2025a) An Yang, Anfeng Li, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chang Gao, Chengen Huang, Chenxu Lv, et al. 2025a. Qwen3 technical report. arXiv preprint arXiv:2505.09388 (2025).
- Yang et al. (2025b) An Yang, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoyan Huang, Jiandong Jiang, Jianhong Tu, Jianwei Zhang, Jingren Zhou, et al. 2025b. Qwen2. 5-1M Technical Report. arXiv preprint arXiv:2501.15383 (2025).
- Yao et al. (2024) Yuan Yao, Tianyu Yu, Ao Zhang, Chongyi Wang, Junbo Cui, Hongji Zhu, Tianchi Cai, Haoyu Li, Weilin Zhao, Zhihui He, et al. 2024. Minicpm-v: A gpt-4v level mllm on your phone. arXiv preprint arXiv:2408.01800 (2024).
- Ye et al. (2023) Qinghao Ye, Haiyang Xu, Guohai Xu, Jiabo Ye, Ming Yan, Yiyang Zhou, Junyang Wang, Anwen Hu, Pengcheng Shi, Yaya Shi, et al. 2023. mplug-owl: Modularization empowers large language models with multimodality. arXiv preprint arXiv:2304.14178 (2023).
- Yu et al. (2023) Weihao Yu, Zhengyuan Yang, Linjie Li, Jianfeng Wang, Kevin Lin, Zicheng Liu, Xinchao Wang, and Lijuan Wang. 2023. Mm-vet: Evaluating large multimodal models for integrated capabilities. arXiv preprint arXiv:2308.02490 (2023).
- Yue et al. (2024) Xiang Yue, Yuansheng Ni, Kai Zhang, Tianyu Zheng, Ruoqi Liu, Ge Zhang, Samuel Stevens, Dongfu Jiang, Weiming Ren, Yuxuan Sun, et al. 2024. Mmmu: A massive multi-discipline multimodal understanding and reasoning benchmark for expert agi. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 9556–9567.
- Zhang et al. (2024) Yi-Fan Zhang, Huanyu Zhang, Haochen Tian, Chaoyou Fu, Shuangqing Zhang, Junfei Wu, Feng Li, Kun Wang, Qingsong Wen, Zhang Zhang, Liang Wang, Rong Jin, and Tieniu Tan. 2024. MME-RealWorld: Could Your Multimodal LLM Challenge High-Resolution Real-World Scenarios that are Difficult for Humans? arXiv preprint arXiv:2408.13257 (2024).
- Zhang et al. (2025) Zicheng Zhang, Xiangyu Zhao, Xinyu Fang, Chunyi Li, Xiaohong Liu, Xiongkuo Min, Haodong Duan, Kai Chen, and Guangtao Zhai. 2025. Redundancy Principles for MLLMs Benchmarks. arXiv:2501.13953 [cs.CL]
- Zhu et al. (2023) Deyao Zhu, Jun Chen, Xiaoqian Shen, Xiang Li, and Mohamed Elhoseiny. 2023. Minigpt-4: Enhancing vision-language understanding with advanced large language models. arXiv preprint arXiv:2304.10592 (2023).
- Zhu et al. (2025) Jinguo Zhu, Weiyun Wang, Zhe Chen, Zhaoyang Liu, Shenglong Ye, Lixin Gu, Yuchen Duan, Hao Tian, Weijie Su, Jie Shao, et al. 2025. InternVL3: Exploring Advanced Training and Test-Time Recipes for Open-Source Multimodal Models. arXiv preprint arXiv:2504.10479 (2025).